# Uncovering the Hidden Threat of Text Watermarking from Users with Cross-Lingual Knowledge

### v1

@@ 1.1 Perplexity (PPL)
where $\exp$ represents the exponential function. $N$ is the total number of words in the text. $p(w_{i})$ is the probability assigned to the word $w_{i}$ by the model. The negative log likelihood, $-\log p(w_{i})$, is referred to the Cross-Entropy loss for word $w_{i}$.
@@ 2.1 Complete Self-BLEU Results
3 : Complete SelfBleu results with the Adjusted Diversity (AD) metric for all watermark methods. For KGW, Unigram, and XSIR, we fix $\gamma=0.5$ and vary $\delta=[2.0,5.0,10.0]$
@@ 3.1 Complete Soft-Win Results
In Table  4, we show more results across different hyper-parameter settings for KGW, Unigram and XSIR. The soft-win rates drastically decreases with lower $\gamma$ values and higher $\delta$ values.
@@
B Detailed Detection Results
@@ 1.2 Watermark Detection with Multiple $\gamma$ Ratios for KGW and Unigram
KGW  Kirchenbauer et al. (2023a) as shown in  3(a) performs the best among all other watermark methods, while Unigram as shown in figure  3(b) performs the worst. XSIR and EXP as shown in figures  3(c) and  3(d) respectively show higher TPRs compared to KGW and Unigram. In all watermark methods, English language shows stable AUC curve in all FPRs in all watermark methods. For KGW and Unigram, increasing the value of $\gamma$ and decreasing the value of $\delta$ has adverse effect on the TPR characteristic, especially for languages other than English (see subsequent sections for roc curves with zoomed-in rates).
Kirchenbauer et al. (2023a) has investigated the tight relationship of $\gamma$ and $\delta$ in watermark strength by creating a lower bound on $\gamma$ with the spike entropy in the picture. However, we believe more investigation is needed in cross-lingual manner. For example, beside the effect of the sampling method (greedy or multinomial), what could be done to incorporate the role of the tokenization of different languages in the $\gamma$ lower bound analysis? By doing a simple analysis of the results with $\gamma=0.9$ and $\delta=0.2$, we found that only an average of $2$ greenlist tokens are missing to raise the z-score from, say $3.52$ to $4$, with a z-score of threshold $4$. As shown in Figures 7(a) and 7(b), the z-score threshold performs adequately for both methods in English but struggles with other languages. At FPR 0%, KGW experiences the most significant TPR drop in Arabic and Indonesian, while Unigram reach worst TPR primarily in Chinese and Indonesian. This suggests that the current z-score setting may not be optimal for all languages, particularly for this $\gamma$ configuration. Further investigation into language-specific z-score adjustments is left for future work.
While this is more evident in KGW and Unigram, XSIR as shown in Figure 8(a) has relatively minor effects for smaller values of $\delta$. This is because XSIR logically uses $50\%$ of the total vocabulary, which is high compared to a smaller $\delta$ value such as $2.0$.
For the Exponential Minimum Sampling or EXP as shown in Figure  8(b), all languages are closely stable in terms of the curve characteristics.
@@
6 : Performance Metrics for Multiple Languages and Watermarking Methods calculated by following the test statistics score. The results show the effect of varying $\delta$ values with a fixed of $\gamma=0.5$ for KGW, Unigram, and XSIR. Using a threshold of $0.2$ for XSIR resulted in higher FPRs Specifically for the English language. (see note below).
(5)(5)footnotetext:
XSIR logically employs 50-50 split ratio. It’s not a hyper-parameter that can be changed directly.
(a) Translation-¿Paraphrase Attacks
(b) Translation-¿Paraphrase-¿Translation Attacks

### v2

@@ 1.1 Perplexity (PPL)
where $\exp$ represents the exponential function. $N$ is the total number of words in the text. $p(w_{i})$ is the probability assigned to the word $w_{i}$ by the model. The negative log likelihood, $\log p(w_{i})$, is referred to the Cross-Entropy loss for word $w_{i}$.
@@ 2.1 Complete Self-BLEU Results
3.1 Sensitivity Analysis for $w$ choice in AD metric
The choice of $w$ in our paper is based on empirical results. In Table 4 we show more resuts with varying the $w$ parameters to be in $[0.1,0.3,0.7]$. After we generate the outputs for a specific language, and after we noticed unexplainable numbers for some results (for example very low Self-Bleu scores), we revise the outputs to check if something is off with the data. Self-blue (SB) results are sometimes <0.05<0.05 which is good in terms of diversity, yet the quality of the output is not as good due to mixture of characters from different languages or gibberish chars, hence lower values of SB scores. However, when the text is generated correctly, we want to catch any repetition in the text due to watermarking. The result of the Judge could still help here since we only utilize its explanation about the coherency of the text, and not its final verdict as to which text is better to avoid judger choice biases. For this reason, we preferred higher weight when we used the coherency scores. Additionally, larger $w$ weights will attribute more weight to SB which is mere n-gram test. Therefore, any non-repeated but incorrect generations will not be explained in the SB scores.
3 : Complete SelfBleu results with the Adjusted Diversity (AD) metric with $w=0.3$ for all watermark methods. For KGW, Unigram, and XSIR, we fix $\gamma=0.5$ and vary $\delta=[2.0,5.0,10.0]$
@@ 3.1 Sensitivity Analysis for $w$ choice in AD metric
4 : Self-BLEU / Adjusted Diversity for All Methods, Languages, and Weights in $w=[0.1,0.3,0.7]$ for CohereAI model. The $w$ weight is used in equation 1 where smaller values indicate larger weighing from judger coherency scores. Smaller values like 0.1 and 0.3 seems to reflect closer AD scores for the watermarked text in comparison to a large one such as 0.7.
Method Lang $w$ Watermarked SB / AD Unwatermarked SB / AD
KGW En 0.1 0.16 / 0.44 0.16 / 0.40
0.3 0.16 / 0.38 0.16 / 0.34
0.7 0.16 / 0.25 0.16 / 0.24
Ar 0.1 0.11 / 0.43 0.12 / 0.42
0.3 0.11 / 0.36 0.12 / 0.35
0.7 0.11 / 0.22 0.12 / 0.22
Zh 0.1 0.04 / 0.55 0.04 / 0.51
0.3 0.04 / 0.44 0.04 / 0.40
0.7 0.04 / 0.21 0.04 / 0.19
Id 0.1 0.10 / 0.52 0.10 / 0.46
0.3 0.10 / 0.43 0.10 / 0.38
0.7 0.10 / 0.24 0.10 / 0.22
Unigram En 0.1 0.23 / 0.45 0.17 / 0.40
0.3 0.23 / 0.40 0.17 / 0.35
0.7 0.23 / 0.30 0.17 / 0.24
Ar 0.1 0.16 / 0.47 0.12 / 0.42
0.3 0.16 / 0.41 0.12 / 0.35
0.7 0.16 / 0.27 0.12 / 0.22
Zh 0.1 0.02 / 0.60 0.04 / 0.51
0.3 0.02 / 0.47 0.04 / 0.40
0.7 0.02 / 0.21 0.04 / 0.20
Id 0.1 0.12 / 0.55 0.10 / 0.45
0.3 0.12 / 0.46 0.10 / 0.37
0.7 0.12 / 0.27 0.10 / 0.22
XSIR En 0.1 0.20 / 0.57 0.17 / 0.37
0.3 0.20 / 0.49 0.17 / 0.32
0.7 0.20 / 0.32 0.17 / 0.24
Ar 0.1 0.14 / 0.54 0.12 / 0.38
0.3 0.14 / 0.45 0.12 / 0.33
0.7 0.14 / 0.27 0.12 / 0.21
Zh 0.1 0.03 / 0.70 0.04 / 0.48
0.3 0.03 / 0.55 0.04 / 0.38
0.7 0.03 / 0.25 0.04 / 0.19
Id 0.1 0.12 / 0.65 0.11 / 0.44
0.3 0.12 / 0.53 0.11 / 0.37
0.7 0.12 / 0.29 0.11 / 0.22
EXP En 0.1 0.19 / 0.67 0.17 / 0.33
0.3 0.19 / 0.57 0.17 / 0.29
0.7 0.19 / 0.35 0.17 / 0.23
Ar 0.1 0.20 / 0.65 0.12 / 0.37
0.3 0.20 / 0.55 0.12 / 0.31
0.7 0.20 / 0.35 0.12 / 0.21
Zh 0.1 0.13 / 0.74 0.04 / 0.48
0.3 0.13 / 0.61 0.04 / 0.38
0.7 0.13 / 0.34 0.04 / 0.19
Id 0.1 0.21 / 0.67 0.10 / 0.43
0.3 0.21 / 0.57 0.10 / 0.36
0.7 0.21 / 0.36 0.10 / 0.21
@@ 4.1 Complete Soft-Win Results
In Table 5, we show more results across different hyper-parameter settings for KGW, Unigram and XSIR. The soft-win rates drastically decreases with lower $\gamma$ values and higher $\delta$ values. This is explainable since larger $\delta$ values increases the magnitude of watermark signal in which the next-token generation is greatly affected.
@@
5.1 More Quality Results for Turkish and Hindi Languages
In Figure 7, we show more coherency difference results for low-resource languages such as Turkish and Hindi using the CohereAI model.
7 : GPT-judge coherency criterion results. We compute the average of watermarked and unwatermarked scores for $500$ generations. Here we only use Cohere model for more languages like Turkish and Hindi.
B More Detection Results
(a) KGW
(b) Unigram
(c) XSIR
(d) EXP
@@ B More Detection Results
(a) Translation-¿Paraphrase Attacks
(b) Translation-¿Paraphrase-¿Translation Attacks
@@ 1.2 Watermark Detection with Multiple $\gamma$ Ratios for KGW and Unigram
KGW  Kirchenbauer et al. (2023a) as shown in  8(a) performs the best among all other watermark methods, while Unigram as shown in figure  8(b) performs the worst. XSIR and EXP as shown in figures  8(c) and  8(d) respectively show higher TPRs compared to KGW and Unigram. In all watermark methods, English language shows stable AUC curve in all FPRs in all watermark methods. For KGW and Unigram, increasing the value of $\gamma$ and decreasing the value of $\delta$ has adverse effect on the TPR characteristic, especially for languages other than English (see subsequent sections for roc curves with zoomed-in rates).
Kirchenbauer et al. (2023a) has investigated the tight relationship of $\gamma$ and $\delta$ in watermark strength by creating a lower bound on $\gamma$ with the spike entropy in the picture. However, we believe more investigation is needed in cross-lingual manner. For example, beside the effect of the sampling method (greedy or multinomial), what could be done to incorporate the role of the tokenization of different languages in the $\gamma$ lower bound analysis? By doing a simple analysis of the results with $\gamma=0.9$ and $\delta=0.2$, we found that only an average of $2$ greenlist tokens are missing to raise the z-score from, say $3.52$ to $4$, with a z-score of threshold $4$. As shown in Figures 9(a) and 9(b), the z-score threshold performs adequately for both methods in English but struggles with other languages. KGW experiences TPR drop in Arabic and Indonesian, while Unigram’s TPR drops primarily in Chinese and Indonesian with low FPR rates. XSIR as shown in Figure 10(a) has relatively minor effects for smaller values of $\delta$. This is because XSIR logically uses $50\%$ of the total vocabulary, which is high compared to a smaller $\delta$ value such as $2.0$. Finally, EXP, as shown in Figure  10(b), shows stable results for all languages in terms of the curve characteristics.
We further show results in Tables  6,  8 for C4 dataset, and Table 7 for LFQA dataset. In these tables we show the results when fixing the $z$-threshold to 4.0, 5.0 for KGW, Unigram, XSIR, and $p$-value to 10e510e-5 and 10e410e-4 for EXP method. The same conclusion can be drawn from using larger $\gamma$ values with smaller $\delta$ values. Using larger values of the threshold makes detection harder for those $\gamma$ settings as shown in Table  6 for all languages. The effect of detection drop is minimized when LFQA instruction-following task is used as shown in Table 7 with exception of Unigram method where TPRs drops are clear for fixed $z$-threshold.
@@
7 : Performance Metrics for Multiple Languages and Watermarking Methods calculated by following the test statistics score with fixed $z=4.0$ and automatic $z$ threshold for LFQA dataset. Here we employ $\gamma=0.5$ and $\delta=2.0$ for KGW, Unigram and XSIR.
Method Language Metrics ($z=4.0$) Automatic $z$ thresholds
TPR TNR FPR FNR $0.1\% TPR@FPR$$1\%$
KGW English 0.854 1.000 0.000 0.146 0.958 0.984
Arabic 0.834 1.000 0.000 0.166 0.936 0.986
Chinese 0.942 0.998 0.002 0.058 0.918 0.998
Indonesian 0.972 1.000 0.000 0.028 0.990 1.000
Unigram English 0.490 1.000 0.000 0.510 0.626 0.944
Arabic 0.568 1.000 0.000 0.432 0.648 0.938
Chinese 0.956 0.988 0.012 0.044 0.876 0.954
Indonesian 0.616 1.000 0.000 0.384 0.894 0.980
EXP English 0.980 1.000 0.000 0.020 0.990 0.992
Arabic 0.996 1.000 0.000 0.004 1.000 1.000
Chinese 1.000 1.000 0.000 0.000 1.000 1.000
Indonesian 1.000 1.000 0.000 0.000 1.000 1.000
XSIR English 0.958 1.000 0.000 0.042 0.964 0.992
Arabic 0.994 0.920 0.080 0.006 0.866 0.976
Chinese 0.996 0.982 0.018 0.004 0.970 0.994
Indonesian 1.000 0.994 0.006 0.000 0.994 1.000
8 : Performance Metrics for Multiple Languages and Watermarking Methods calculated by following the test statistics score for C4 dataset. The results show the effect of varying $\delta$ values with a fixed of $\gamma=0.5$ for KGW, Unigram, and XSIR. Using a threshold of $0.2$ for XSIR resulted in higher FPRs Specifically for the English language.
@@ 2.2 Watermark Detection After Attacks for Syntactical Methods
C GPT-Judger Fairness Experiments
We perform two experiments in an attempt to elicit any possible biases toward a specific language over the other. First, we focus on the languages we study which are English, Arabic, Chinese and Indonesian. We also add closely related languages such as Persian or Farsi (close to Arabic), German (close to English), and Japanese (close to Chinese). To prepare our data with ground truth values, we use natural English examples from the C4 dataset then we translate these examples to all other languages using a powerful GPT model such as GPT-3.5-Turbo. The objective here is to ensure that we have the same text but in different languages to ascertain any biases toward a specific language. In all of our judging experiments we randomized the order in which the texts are fed to the LLM and use two runs under different seeds to allow for a more rigor investigation.
Judging Texts from Different Languages. To further analyze the judger’s result, we conduct a fairness experiment we call ”Translation Experiment” in which we investigate whether GPT-Judger shows bias toward a specific language when given the same text in multiple languages. We use the same judge we used for our main experiment in the main content of the paper, which means we used the same LLM and system prompt with slight adjustment to the prompt to account for multiple options instead of only two options (unwatermarked and watermarked texts.) Table 9 shows the result of judging the same text in multiple languages. The hard win column reflects the judge final verdict for the language of the winning text. the first-last column reflects the percentage of the winning text when it is placed in the first or the last option in the prompt to investigate position bias if any  (Pezeshkpour and Hruschka, 2023). As can be seen from the table, we see $6.5\%$ for both English and Chinese, yet their hard win rates are largely different. Therefore, it is not clear whether the order affects the final verdict of the judge in this settings. However, it’s clear that some languages are preferred over other despite ensuring that the texts are mere translations of each other, and therefore should have more TIE percentages (5)(5)(5)We instruct the judge to clearly return TIE as the final verdict if the text quality of all languages are equally good.. For example, German language is the most preferred one, followed by Arabic, and finally English, while Chinese, Persian, and Indonesian languages receives the lowest scores. Persian language, despite the fact that it originates from the same family as Arabic, receives the lowest scores. Therefore, we hypothesize that there is a bias toward languages that the judge is most proficient at or more familiar with.
9 : Translation Verdict Percentages over two runs with different seeds. Each run contains 100 samples from 500 translation examples for a total of 200 examples. The first-last column reflects the percentage of the winning text when it is placed in the first or the last option in the prompt to investigate position bias if any.
Language hard win (%) first-last (%)
EN 19.50 $\pm$ 4.95 6.50
JA 14.50 $\pm$ 3.54 13.00
FA 2.00 $\pm$ 0.00 2.00
AR 23.00 $\pm$ 5.66 13.00
ZH 7.00 $\pm$ 1.41 6.50
DE 30.50 $\pm$ 3.54 11.00
ID 3.50 $\pm$ 0.71 3.50
Total 100 55.5
10 : Paraphrase Verdict Percentages over two runs with different seeds. Each run contains 100 samples from 500 paraphrased examples from the original non-perturbed text. The Perturbed Text column is the paraphrased version of the Natural Text.
Language Perturbed Text (%) Natural Text (%) TIE (%) Model Failure (%)
ZH 30.00 $\pm$ 1.41 26.00 $\pm$ 0.00 44.00 $\pm$ 1.41 0.00
FA 39.50 $\pm$ 3.54 20.00 $\pm$ 4.24 40.50 $\pm$ 0.71 0.00
JA 34.00 $\pm$ 7.07 28.50 $\pm$ 0.71 37.50 $\pm$ 6.36 0.00
AR 40.50 $\pm$ 4.95 30.50 $\pm$ 6.36 29.00 $\pm$ 1.41 0.00
DE 47.50 $\pm$ 6.36 28.50 $\pm$ 2.12 23.00 $\pm$ 8.49 1.00
ID 48.00 $\pm$ 4.24 30.50 $\pm$ 7.78 21.50 $\pm$ 3.54 0.00
EN 75.50 $\pm$ 9.19 24.00 $\pm$ 9.90 0.50 $\pm$ 0.71 0.00
Judging Texts from the Same Language In this experiment which we call ”Paraphrase Experiment”, we used the same judge as the one used watermarking experiments. Our natural text in this case represent the original translated text (for non-English) from the translation experiment. In this experiment, the two texts (natural and perturbed) are essentially the same. In other words, we use GPT-3.5-turbo to paraphrase the original text and we call the resultant text the perturbed text. We create 500 examples in this manner and we used two seeds to sample different 100 examples for our fairness assessments. Table  10 shows the results from using GPT-Judger with these modified texts to assess their quality. Comparing the paraphrase results with that of the translation, we see a pattern that confirms the translation experiments. We noticed that the languages that received lower scores in translation experiment are at the top of the paraphrase tables, and those that received higher scores are at the bottom in terms of the TIE scores. However, there is an exception for the Indonesian language, which appears to have lower TIE scores. We could argue here that with the model being able to say two texts are equal in quality is not because it is able to assert that the quality of the two texts are equally good, but rather because it might not be proficient in these languages (as seen in Table 9) that it preferred to choose a TIE. For Indonesian language, token bias as investigated by Zheng et al. can play a role since the alphabets used in this language are the same as the English alphabets. Another important observation is that the perturbed text is favored more by the LLM judger in all the languages as shown in the second column of the table, whereas natural text is less favored by the judger despite not being gone through a lot of modification. This confirms studies that indicate LLMs favoring texts generated by their own model variants  (Ye et al., 2024) since our perturbed text has gone through perturbations of the same model variant. Therefore, it has become easier for the GPT-judge to select as the winning text. However, in our main watermarking experiments, our perturbed texts (watermarked texts and non-English language texts) don’t have such perturbation footprint, which can put aside the bias toward model-own generations from our main experiments.